OpenStudy (anonymous):

Write the Equation in Slope-Intercept form. What are the slope and y intercept? -2x+4y= -10

OpenStudy (solomonzelman):

slope intercept form is \[y=mx+b\]m - slope y - y intercept. \[-2x+4y=-10\]divide everything in the equation by 4,\[\color{black}{ \frac{1}{2} }x+ y=-\color{black}{ \frac{5}{2} } \]set it to y=mx+b form,\[\color{black}{ - \frac{1}{2} x } +y\color{blue}{ - \frac{1}{2} x } =\color{black}{ - \frac{5}{2} } \color{blue}{ - \frac{1}{2} x } \]\[y=\color{black}{ - \frac{1}{2} x } -\color{black}{ \frac{5}{2} } \]

OpenStudy (solomonzelman):

-1/2 = slope (or "m" ) -5/2 = y-intercept.